OpenStudy (anonymous):

Help!!! Explain what factored form a quadratic function helps you to graph and give and example. Please help!

OpenStudy (cwrw238):

the binomial and vertex form

OpenStudy (anonymous):

Ohhh okay (:

OpenStudy (cwrw238):

an example: y = x^2 + 4x - 5 factor: y = (x + 5)(x - 1) - so the zeros (x-intercepts) are -5 and 1 vertex form y = (x + 2)^2 - 9 - this gives you the coordinates of the minimum value , in this case its (-2,6)

OpenStudy (cwrw238):